OpenStudy (anonymous):

classify the graph of the equation as a circle, a parabola, an ellipse, or a hyperbola. Rewrite each equation in standard form. please help!

OpenStudy (anonymous):

10x^2 +3y^2 + 9x + 10y -10 =0

OpenStudy (thatonegirl_):

Do you still need help?

OpenStudy (anonymous):

yeah

OpenStudy (thatonegirl_):

Well anyways, here are the rules for identifying between all of those. A circle will have the SAME coefficient in front of the x^2 and the y^2, and they will also have a + between them. A parabola will only have one of the terms (either the x or y) squared. An ellipse will have an x^2 and y^2 but different coefficients, and they will have a plus sign in between them. And a hyperbola will be the same as an ellipse except with a minus sign. This should be helpful: http://www.dummies.com/how-to/content/how-to-identify-the-four-conic-sections-in-equatio.html

OpenStudy (anonymous):

i need help writing it in standard form @Thatonegirl_

OpenStudy (thatonegirl_):

Oh ok. Circle is (x – h)^2 + (y – k)^2 = r^2

OpenStudy (thatonegirl_):

An ellipse is: \[\frac{ (x-h)^2 }{ a^2 }+\frac{ (y-k)^2 }{ b^2}=1\]

OpenStudy (thatonegirl_):

and the a and b can flip

OpenStudy (anonymous):

i am talking about getting the equation to standard form

OpenStudy (thatonegirl_):

do u have an equation ?

OpenStudy (anonymous):

not the standard form for each conic shape

OpenStudy (anonymous):

yeah its under my question

OpenStudy (anonymous):

second line

OpenStudy (thatonegirl_):

okay move all the x's together and the y's and add the -10 to the other side.

OpenStudy (thatonegirl_):

After u do that, complete the square for the x's and y's

OpenStudy (anonymous):

alrigth

OpenStudy (thatonegirl_):

Factor each and then divide by the 10 you added to the other side.

OpenStudy (anonymous):

i get 10x^2 + 9x + 81/4 +3y^2 +10y +25 = 221/4

OpenStudy (anonymous):

is it right

OpenStudy (thatonegirl_):

Oh sorry you are supposed to factor the 10 and the 3 from each before you complete the square

OpenStudy (thatonegirl_):

I'll go through it, stop me if you have a question.

OpenStudy (thatonegirl_):

\[10x^2 +3y^2 + 9x + 10y -10 =0\]

OpenStudy (anonymous):

okay

OpenStudy (thatonegirl_):

\[10x^2+9x^2+3y^2+10y=10\]

OpenStudy (thatonegirl_):

\[10(x^2+\frac{ 9 }{ 10 }+ )+3(y^2+\frac{ 10 }{ 3 })=10\]

OpenStudy (thatonegirl_):

\[10(x^2+\frac{ 9 }{ 10 }+\frac{ 81 }{ 400 })+3(y^2+\frac{ 10 }{ 3 }+\frac{ 25 }{ 9 })=10+\frac{ 25 }{ 3 }+\frac{ 81 }{ 40}\]

OpenStudy (anonymous):

do you have time to help me with one more?

OpenStudy (thatonegirl_):

Get all that?

OpenStudy (anonymous):

yeah i see how you did it now

OpenStudy (anonymous):

this other is different though it doesnt have a y^2 just an x^2

OpenStudy (thatonegirl_):

What?

OpenStudy (anonymous):

x^2+4y+7x-10 =0

OpenStudy (thatonegirl_):

Are we doing the equation you posted?

OpenStudy (thatonegirl_):

Oh so that's going to be a parabola, correct?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

like this is what i have so far x^2+7x+49/4 = -4y + 89/4

OpenStudy (anonymous):

this is for the second problem

OpenStudy (thatonegirl_):

Well what I do is move everything except the x^2 and x to the other side. Then factor out the coefficient.

OpenStudy (anonymous):

do i factor out the x ------->x (x+7) = -2(2y-5)? @thatonegirl_